# Fluctuations for non-Hermitian dynamics

### v1

@@ 4 Independence of singular values
${\rm d}m_{s}(z_{s})=-$ $\frac{1}{2\sqrt{2}\normalcolor N^{3/2}}\sum_{i}\frac{{\rm d}B_{i}(s)}{(x_{i}(s)-z_{s})^{2}}+(m_{s}(z_{s})-\widetilde{m}_{s}(z_{s}))\partial_{z}m_{s}(z_{s}){\rm d}s+\frac{1}{N^{2}}\sum_{i}\frac{\operatorname{O}(1){\rm d}s}{(x_{i}(s)-z_{s})^{3}}.$ (4.44)
@@
$\int_{0}^{s\wedge\sigma}\frac{1}{N^{3/2}}\sum_{i=1}^{N}\frac{{\rm d}B_{i}(\tau)}{|z_{\tau}(u)-x_{i}(\tau)|^{2}}\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{s\wedge\sigma}(z_{s\wedge\sigma})]}{N\eta_{s\wedge\sigma}}},$ (4.46)
$\int_{0}^{s\wedge\sigma}\frac{1}{N^{2}}\sum_{i=1}^{N}\frac{1}{|x_{i}(\tau)-z_{\tau}(u)|^{3}}{\rm d}\tau\lesssim\frac{1}{N\eta_{s\wedge\sigma}(u)}.$ (4.47)
@@
$\frac{1}{N^{2}}\int_{0}^{s}\sum_{ij}\frac{{\rm d}\tau}{|x_{i}(\tau)-z_{\tau}|^{2}|x_{j}(\tau)-z_{\tau}|^{2}}=\int_{0}^{s}\frac{{\rm Im}[m_{\tau}(z_{\tau})]^{2}}{\eta_{\tau}^{2}}{\rm d}\tau\leqslant\int_{0}^{s}\frac{{\rm Im}[\widetilde{m}_{\tau}(z_{\tau})]^{2}}{\eta_{\tau}^{2}}{\rm d}\tau\lesssim\frac{{\rm Im}[\widetilde{m}_{s}(z_{s})]}{\eta_{s}}$ (4.49)
where we successively used (4.45) and $-{\rm Im}[\widetilde{m}_{\tau}(z_{\tau})]=\partial_{\tau}\eta_{\tau}$. With (4.33) and (4.49) have proved that for some $c_{1}>0$, with probability $1-c_{1}^{-1}e^{-c_{1}\varphi^{1/10}}$ we have
$\sup_{0\leqslant{\tau}\leqslant t^{(k)}}\left|\int_{0}^{{\tau}\wedge\sigma}\frac{1}{N^{3/2}}\sum_{i=1}^{N}\frac{{\rm d}B_{i}({\tau})}{|z_{\tau}-x_{i}({\tau})|^{2}}\right|\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{t^{(k)}\wedge\sigma}(z_{t^{(k)}\wedge\sigma})]}{N\eta_{t^{(k)}\wedge\sigma}}}.$ (4.50)
@@
$\begin{split}\left|\int_{0}^{s\wedge\sigma}\frac{1}{N^{3/2}}\sum_{i=1}^{N}\frac{{\rm d}B_{i}(\tau)}{|z_{\tau}-x_{i}(\tau)|^{2}}\right|\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{t^{(k)}\wedge\sigma}(z_{t^{(k)}\wedge\sigma})]}{N\eta_{t^{(k)}\wedge\sigma}}}\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{s\wedge\sigma}(z_{s\wedge\sigma})]}{N\eta_{s\wedge\sigma}}}.\end{split}$
@@
$\int_{0}^{t\wedge\sigma}\frac{\operatorname{O}(1)}{N^{2}}\sum_{i=1}^{N}\frac{{\rm d}s}{|x_{i}(s)-z_{s}|^{3}}\leqslant\int_{0}^{t\wedge\sigma}\frac{\operatorname{O}(1)}{N}\frac{{\rm Im}[m_{s}(z_{s})]{\rm d}s}{\eta_{s}^{2}}\leqslant\int_{0}^{t\wedge\sigma}\frac{\operatorname{O}(1)}{N}\frac{{\rm Im}[\widetilde{m}_{s}(z_{s})]{\rm d}s}{\eta_{s}^{2}}\lesssim\frac{1}{N\eta_{t\wedge\sigma}}.$
The first inequality relies on $|x_{i}(s)-z_{s}|\geqslant\eta_{s}$, and last inequality uses $-{\rm Im}[\widetilde{m}_{s}(z_{s})]=\partial_{s}\eta_{s}$, see (4.38). ∎
@@
$\left|\partial_{z}m_{v}(z_{v}(u))\right|\leqslant\partial_{z}\widetilde{m}_{v}(z_{v}(u))+\frac{1}{\eta_{v}}\sqrt{\varphi{\rm Im}[\widetilde{m}_{v}(z_{v})]\sqrt{\frac{{\rm Im}[\widetilde{m}_{v}(z_{v})]}{N\eta_{v}}}}=\partial_{z}\widetilde{m}_{v}(z_{v}(u))+\varphi^{\frac{1}{2}}\frac{{\rm Im}[\widetilde{m}_{v}(z_{v})]^{3/4}}{\eta_{v}(N\eta_{v})^{1/4}}.$ (4.52)
@@
Before proving Proposition 4.6, we explain that the obtained result is optimal, up to subpolynomial (in $N$) errors. Without loss of generality we assume that $\widetilde{\rho}_{t}(0)=\widetilde{\rho}^{\prime}_{t}(0)$, and apply the local law estimate (4.9) to obtain the best possible bound on the inital condition of Equation (4.17): $|v_{i}(0)|=|s_{i}(0)-s^{\prime}_{i}(0)|\lesssim\frac{\sqrt{i}}{N}$ and $v_{-i}(0)=-v_{i}(0)$. As the expected continuous limit of the propagator ${\mathcal{U}}$ of (4.17) is $p_{s,t}(x,y)=\frac{t-s}{(t-s)^{2}+(x-y)^{2}}$ (see [BouErdYauYin2016]), a perfect homogenization of (4.17) would give
$|v_{i}(\alpha,t)|\lesssim\frac{1}{N}\int\frac{t}{t^{2}+(x-\frac{i}{N})^{2}}\cdot\operatorname{sgn}(x)(Nx)^{\frac{1}{2}}{\rm d}x\lesssim\frac{i}{N}\cdot\frac{1}{(\max(i,Nt))^{\frac{1}{2}}}.$
@@
Remark 4.17.
If we assume that the initial condition is rigid instead of the weak local law (4.9), i.e. if $|v_{i}(0)|=|s_{i}(0)-s^{\prime}_{i}(0)|\lesssim\frac{1}{N}$, then the above heuristic gives $|s_{i}(t)-s^{\prime}_{i}(t)|\lesssim\frac{i}{N}\cdot\frac{1}{\max(i,Nt)}$. The methods developed below likely apply to cover such stronger assumptions and results, but we do not pursue this direction in this article.
Thanks to Assumption 4.3, the two measures $\widetilde{\mu}_{0},\widetilde{\mu}^{\prime}_{0}$ have bounded densities $\widetilde{\rho}_{0}(x),\widetilde{\rho}^{\prime}_{0}(x)$, and their densities have bounded derivatives for $|x|\leqslant G$. We can first do a rescaling, to make that the two densities are the same at 00: $\widetilde{\rho}_{0}(0)=\widetilde{\rho}^{\prime}_{0}(0)=A$. Since $t\ll 1$, the free convolution density also satisfies $\widetilde{\rho}_{t}(0)=\widetilde{\rho}^{\prime}_{t}(0)=A+\operatorname{O}(t)$. Thus Proposition 4.6 follows if we can show that
$|s_{i}(t)-s_{i}^{\prime}(t)|\leqslant N^{\varepsilon}\cdot\frac{|i|}{N}\cdot\left(\frac{1}{\sqrt{Nt}}+\max\left(\frac{|i|}{N},t\right)\right),\quad|i|\leqslant N.$
We now start the proof of Proposition 4.6. Assumption 4.3 implies that $\widetilde{\rho}_{0}(x),\widetilde{\rho}^{\prime}_{0}(x)$ have bounded derivatives for $|x|\leqslant G$, and it follows
@@
$\frac{i}{N}=\int_{0}^{\gamma_{i}}\widetilde{\rho}_{0}(x){\rm d}x,\quad|i|\geqslant 1,$
Then for $|i|\leqslant c_{G}N$, (4.56) implies that
$\gamma_{i}=\frac{i}{AN}+\operatorname{O}\left(\frac{i^{2}}{N^{2}}\right).$
By a standard argument (see [Lazlo2009, Corollary 4.2]), $(\nu,G)$-regularity in the sense of Assumption 4.3 implies that
@@
for any interval $I\subset[-c_{G},c_{G}]$ with $|I|\geqslant N^{-1+\nu}$. Since $\mu_{0}$ and $\widetilde{\mu}_{0}$ are both symmetric around origin, we conclude that for any $|i|\leqslant cN$
$|s_{i}(0)-\gamma_{i}(0)|\lesssim\frac{i^{1/2}+N^{\nu}}{N}+\frac{i^{2}}{N^{2}}.$
@@
$|s^{\prime}_{i}(0)-\gamma_{i}(0)|\lesssim\frac{i^{1/2}+N^{\nu}}{N}+\frac{i^{2}}{N^{2}},\qquad\quad|s_{i}(0)-s^{\prime}_{i}(0)|\lesssim\frac{i^{1/2}+N^{\nu}}{N}+\frac{i^{2}}{N^{2}}.$ (4.58)
@@
where $x_{i}(0,\alpha)=\alpha s_{i}+(1-\alpha)s^{\prime}_{i},\quad 0\leqslant\alpha\leqslant 1,$ and the Brownian motions satisfies ${\rm d}\langle B_{i}\rangle/{\rm d}t\leqslant 1$. We consider the corresponding tangential equation (4.17) satisfied by $v_{i}(t)=v_{i}(\alpha,t)=\frac{{\rm d}}{{\rm d}\alpha}x_{i}(\alpha,t)$
@@
${\rm d}f_{t}=m_{t}(z)\partial_{z}f_{t}{\rm d}t+\frac{1}{2N}\sum_{|k|\leqslant N}\frac{v_{k}(t)({\rm d}\langle B_{k}\rangle_{t}-{\rm d}t)}{(x_{k}(t)-z)^{3}}-\frac{1}{\sqrt{2N}}\sum_{|k|\leqslant N}\frac{v_{k}(t)}{(z-x_{k}(t))^{2}}{\rm d}B_{k}(t).$
@@
${\rm d}f=\sum_{|k|\leqslant N}\frac{{\rm d}v_{k}}{x_{k}-z}+\sum_{|k|\leqslant N}v_{k}{\rm d}\frac{1}{x_{k}-z}.$ (4.61)
Applying again the Itô formula ${\rm d}(x_{k}-z)^{-1}=-(x_{k}-z)^{-2}{\rm d}x_{k}+\frac{{\rm d}\langle B_{k}\rangle}{2N}(x_{k}-z)^{-3}{\rm d}t$, with (4.59) we naturally decompose the second sum above as (I)+[(II)+(III)]${\rm d}t$ where
${\rm(I)}$ $=-\frac{1}{\sqrt{2N}}\sum_{|k|\leqslant N}\frac{v_{k}}{(z-x_{k})^{2}}{\rm d}B_{k},$
@@
${\rm(III)}$ $=\frac{1}{2N}\sum_{|k|\leqslant N}\frac{v_{k}{\rm d}\langle B_{k}\rangle}{(x_{k}-z)^{3}}.$
@@
$\sum_{|k|\leqslant N}\frac{\partial_{t}v_{k}}{x_{k}-z}$ $=\sum_{\ell\neq k}\frac{v_{\ell}-v_{k}}{2N(x_{\ell}-x_{k})^{2}(x_{k}-z)}=\frac{1}{2}\sum_{\ell\neq k}\frac{v_{\ell}-v_{k}}{2N(x_{\ell}-x_{k})^{2}}\left(\frac{1}{x_{k}-z}-\frac{1}{x_{\ell}-z}\right)$
@@
${\rm(II)}+\sum_{|k|\leqslant N}\frac{\partial_{t}v_{k}}{x_{k}-z}=\frac{1}{2N}\sum_{\ell\neq k}\frac{v_{k}}{x_{\ell}-x_{k}}\frac{1}{x_{k}-z}\left(\frac{1}{x_{k}-z}-\frac{1}{x_{\ell}-z}\right)=\frac{1}{2N}\sum_{\ell\neq k}\frac{v_{k}}{(x_{k}-z)^{2}}\frac{1}{x_{\ell}-z}.$
All singularities have disappeared. We obtained ${\rm(II)}+{\rm(III)}+\sum_{|k|\leqslant N}\frac{\partial_{t}v_{k}}{x_{k}-z}=m(z)\partial_{z}f+\frac{1}{2N}\sum_{|k|\leqslant N}\frac{v_{k}(t)({\rm d}\langle B_{k}\rangle_{t}-{\rm d}t)}{(x_{k}(t)-z)^{3}}.$ Summation of the remaining term (I) concludes the proof. ∎
@@
For any $|z|\leqslant G/4$ with $E\geqslant 0$, $\eta:=\operatorname{Im}z\geqslant N^{-1+\nu}$, we have
@@
$\begin{split}|\operatorname{Im}f^{(2)}_{0}(z)|&=\left|\operatorname{Im}\sum_{k\geqslant 1}\left(\frac{v^{(2)}_{k}(0)}{x_{k}(0)-z}-\frac{v^{(2)}_{k}(0)}{-x_{k}(0)-z}\right)\right|\leqslant\eta\sum_{k\geqslant 1}|v^{(2)}_{k}(0)|\left|\frac{1}{|x_{k}-E|^{2}+\eta^{2}}-\frac{1}{|x_{k}(0)+E|^{2}+\eta^{2}}\right|\\ &=4\eta E\sum_{k\geqslant 1}|v^{(2)}_{k}(0)|\frac{x_{k}(0)}{|x_{k}(0)-z|^{2}|x_{k}(0)+z|^{2}}\leqslant 4\eta E\sum_{k\geqslant 1}\frac{k^{2}}{N^{2}}\frac{x_{k}(0)}{|x_{k}(0)-z|^{2}|x_{k}(0)+z|^{2}}.\end{split}$ (4.62)
From (4.57) and the assumption ${\rm Im}z\geqslant N^{-1+\nu}$ for eigenvalues $x_{k}\leqslant G/2$, and the assumption $x_{k}\leqslant C$ for the complementary regime, we have
@@
$|{\rm Im}f_{0}^{(1)}(z)|\leqslant 4\eta E\sum_{k\geqslant 1}\frac{\sqrt{k}}{N}\frac{x_{k}(0)}{|x_{k}(0)-z|^{2}|x_{k}(0)+z|^{2}}\lesssim\sqrt{N}E\int_{0}^{c}\frac{x^{3/2}\eta}{|z-x|^{2}|z+x|^{2}}{\rm d}x\lesssim E\sqrt{\frac{N}{\max(E,\eta)}},$
@@
$\mathscr{S}=\left\{z=E+\mathrm{i}y:y<E<\frac{G}{10}\right\}.$
@@
$\tau_{0}$ $=\inf\left\{0\leqslant t\leqslant 1\mid\exists z\in[0,G/2]\times[\frac{\varphi^{5}}{N},1]\ {\rm s.t.}\ \left|m_{t}(w)-\widetilde{m}_{t}(w)\right|\geqslant\varphi\sqrt{\frac{{\rm Im}[\widetilde{m}_{0}(w)]}{N{\rm Im}[w]}}\right\},$
$\tau$ $=\min\{\tau_{0},\tau_{\ell,m}:0\leqslant\ell,m\leqslant N^{12}{\mathsf{T}},\varphi^{5}N^{-1}<E_{m}<G/10\},$
@@
$\{\tau={\mathsf{T}}\}\bigcap_{1\leqslant\ell\leqslant N^{12}{\mathsf{T}},1\leqslant m\leqslant N^{12},1\leqslant k\leqslant N}A_{\ell,m,k}\subset\bigcap_{z\in\mathscr{S},0<t<{\mathsf{T}}}\left\{|\operatorname{Im}f_{s}(z_{t-s})-\operatorname{Im}f_{0}(z_{t})|<E\sqrt{\frac{N}{\max\{E,t\}}}+NE\max(E,t)\right\}$ (4.65)

### v2

@@ 4 Independence of singular values
${\rm d}m_{s}(z_{s})=$- $\frac{1}{2\sqrt{2}N^{3/2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}b_{i}(s)}{(x_{i}(s)-z_{s})^{2}}+(m_{s}(z_{s})-\widetilde{m}_{s}(z_{s}))\partial_{z}m_{s}(z_{s}){\rm d}s+\frac{1}{N^{2}}\sum_{1\leqslant|i|\leqslant N}\frac{\operatorname{O}(1){\rm d}s}{(x_{i}(s)-z_{s})^{3}}.$ (4.49)
@@
$\int_{0}^{s\wedge\sigma}\frac{1}{N^{3/2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}b_{i}(\tau)}{|z_{\tau}(u)-x_{i}(\tau)|^{2}}\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{s\wedge\sigma}(z_{s\wedge\sigma})]}{N\eta_{s\wedge\sigma}}},$ (4.51)
$\int_{0}^{s\wedge\sigma}\frac{1}{N^{2}}\sum_{1\leqslant|i|\leqslant N}\frac{1}{|x_{i}(\tau)-z_{\tau}(u)|^{3}}{\rm d}\tau\lesssim\frac{1}{N\eta_{s\wedge\sigma}(u)}.$ (4.52)
@@
$\frac{1}{N^{2}}\int_{0}^{s}\sum_{1\leqslant|i|,|j|\leqslant N}\frac{{\rm d}\tau}{|x_{i}(\tau)-z_{\tau}|^{2}|x_{j}(\tau)-z_{\tau}|^{2}}=\int_{0}^{s}\frac{{\rm Im}[m_{\tau}(z_{\tau})]^{2}}{\eta_{\tau}^{2}}{\rm d}\tau\leqslant\int_{0}^{s}\frac{{\rm Im}[\widetilde{m}_{\tau}(z_{\tau})]^{2}}{\eta_{\tau}^{2}}{\rm d}\tau\lesssim\frac{{\rm Im}[\widetilde{m}_{s}(z_{s})]}{\eta_{s}},$ (4.54)
where we successively used (4.50) and ${\rm Im}[\widetilde{m}_{\tau}(z_{\tau})]=\partial_{\tau}\eta_{\tau}$. With (4.36) and (4.54) have proved that for some $c_{1}>0$, with probability $1-c_{1}^{-1}e^{-c_{1}\varphi^{1/10}}$ we have
$\sup_{0\leqslant{\tau}\leqslant t^{(k)}}\left|\int_{0}^{{\tau}\wedge\sigma}\frac{1}{N^{3/2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}b_{i}({\tau})}{|z_{\tau}-x_{i}({\tau})|^{2}}\right|\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{t^{(k)}\wedge\sigma}(z_{t^{(k)}\wedge\sigma})]}{N\eta_{t^{(k)}\wedge\sigma}}}.$ (4.55)
@@
$\begin{split}\left|\int_{0}^{s\wedge\sigma}\frac{1}{N^{3/2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}b_{i}(\tau)}{|z_{\tau}-x_{i}(\tau)|^{2}}\right|\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{t^{(k)}\wedge\sigma}(z_{t^{(k)}\wedge\sigma})]}{N\eta_{t^{(k)}\wedge\sigma}}}\leqslant\varphi^{1/20}\sqrt{\frac{{\rm Im}[\widetilde{m}_{s\wedge\sigma}(z_{s\wedge\sigma})]}{N\eta_{s\wedge\sigma}}}.\end{split}$
@@
$\int_{0}^{t\wedge\sigma}\frac{\operatorname{O}(1)}{N^{2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}s}{|x_{i}(s)-z_{s}|^{3}}\leqslant\int_{0}^{t\wedge\sigma}\frac{\operatorname{O}(1)}{N}\frac{{\rm Im}[m_{s}(z_{s})]{\rm d}s}{\eta_{s}^{2}}\leqslant\int_{0}^{t\wedge\sigma}\frac{\operatorname{O}(1)}{N}\frac{{\rm Im}[\widetilde{m}_{s}(z_{s})]{\rm d}s}{\eta_{s}^{2}}\lesssim\frac{1}{N\eta_{t\wedge\sigma}}.$
The first inequality relies on $|x_{i}(s)-z_{s}|\geqslant\eta_{s}$, and last inequality uses ${\rm Im}[\widetilde{m}_{s}(z_{s})]=\partial_{s}\eta_{s}$, see (4.41). ∎
@@
$\left|\partial_{z}m_{v}(z_{v}(u))\right|\leqslant\left|\partial_{z}\widetilde{m}_{v}(z_{v}(u))\right|+\frac{1}{\eta_{v}}\sqrt{\varphi{\rm Im}[\widetilde{m}_{v}(z_{v})]\sqrt{\frac{{\rm Im}[\widetilde{m}_{v}(z_{v})]}{N\eta_{v}}}}=\partial_{z}\widetilde{m}_{v}(z_{v}(u))+\varphi^{\frac{1}{2}}\frac{{\rm Im}[\widetilde{m}_{v}(z_{v})]^{3/4}}{\eta_{v}(N\eta_{v})^{1/4}}.$ (4.57)
@@
Before proving Proposition 4.6, we explain why we believe that the obtained result is optimal, up to subpolynomial (in $N$) errors, in the setting of dynamics driven by general correlated Brownian motions and particle locations given by a local law. We denote $\widetilde{\rho}_{t}$ (resp. $\widetilde{\rho}^{\prime}_{t}$) the density of $\widetilde{\mu}_{0}\boxplus\mu_{\rm sc}^{(t)}$ (resp. $\widetilde{\mu}_{0}^{\prime}\boxplus\mu_{\rm sc}^{(t)}$) (see the discussion in Section 4.4). Without loss of generality we assume that $\widetilde{\rho}_{t}(0)=\widetilde{\rho}^{\prime}_{t}(0)$, and apply the local law estimate (4.9) to obtain the best possible bound on the initial condition of Equation (4.20): $|v_{i}(0)|=|s_{i}(0)-s^{\prime}_{i}(0)|\lesssim\frac{\sqrt{i}}{N}$ and $v_{-i}(0)=-v_{i}(0)$. As the expected continuous limit of the propagator $\mathcal{U}$} of (4.20) is $p_{s,t}(x,y)=\frac{t-s}{(t-s)^{2}+(x-y)^{2}}$ (see [BouErdYauYin2016]), a perfect homogenization of (4.20) would give
$|v_{i}(\alpha,t)|\lesssim\frac{1}{N}\int\frac{t}{t^{2}+(x-\frac{i}{N})^{2}}\cdot\operatorname{sgn}(x)(Nx)^{\frac{1}{2}}{\rm d}x\lesssim\frac{i}{N}\cdot\frac{1}{(\max(i,Nt))^{\frac{1}{2}}}.$ (4.61)
@@
Remark 4.19 (Improved estimates under stronger assumptions).
If we assume that the initial condition is rigid instead of the weak local law (4.9), i.e. if $|v_{i}(0)|=|s_{i}(0)-s^{\prime}_{i}(0)|\lesssim\frac{1}{N}$, then the above heuristic gives $|s_{i}(t)-s^{\prime}_{i}(t)|\lesssim\frac{i}{N}\cdot\frac{1}{\max(i,Nt)}$.
In fact, for rigid initial conditions and independent driving Brownian motions, [Wang2022, Theorem 6] proves that this heuristic is correct by adapting the bulk universality proof from [Bou2022]. The assumptions of Proposition 4.6 are more general as they only assume a local law and general correlations for the noise. The methods developed below likely applies to give a simpler proof that $|s_{i}(t)-s^{\prime}_{i}(t)|\lesssim\frac{i}{N}\cdot\frac{1}{\max(i,Nt)}$ under the rigidity and independence assumptions, but we do not pursue this direction in this article.
Thanks to Assumption 4.3, the two measures $\widetilde{\mu}_{0},\widetilde{\mu}^{\prime}_{0}$ have bounded densities $\widetilde{\rho}_{0}(x),\widetilde{\rho}^{\prime}_{0}(x)$, and their densities have bounded derivatives for $|x|\leqslant G$. By our assumption $\widetilde{\rho}_{0}(0)=\widetilde{\rho}^{\prime}_{0}(0)=A$. Since $t\ll 1$, the free convolution density also satisfies $\widetilde{\rho}_{t}(0)=\widetilde{\rho}^{\prime}_{t}(0)=A+\operatorname{O}(t)$. Thus Proposition 4.6 follows if we can show that
$|s_{i}(t)-s_{i}^{\prime}(t)|\leqslant N^{\varepsilon}\cdot\frac{|i|}{N}\cdot\left(\frac{1}{\sqrt{Nt}}+\max\left(\frac{|i|}{N},t\right)\right),\quad 1\leqslant|i|\leqslant N.$
We now start the proof of Proposition 4.6. Assumption 4.3 implies that $\widetilde{\rho}_{0}(x),\widetilde{\rho}^{\prime}_{0}(x)$ have bounded derivatives for $|x|\leqslant R$, and it follows
@@
$\frac{i}{2N}=\int_{0}^{\gamma_{i}}\widetilde{\rho}_{0}(x){\rm d}x,\quad 1\leqslant|i|\leqslant N.$
Then for $|i|\leqslant c_{R}N$, (4.62) implies that
$\gamma_{i}=\frac{i}{2AN}+\operatorname{O}\left(\frac{i^{2}}{N^{2}}\right).$
By a standard argument (see [Lazlo2009, Corollary 4.2]), $(\nu,R)$-regularity in the sense of Assumption 4.3 implies that
@@
for any interval $I\subset[-c_{R},c_{R}]$ with $|I|\geqslant N^{-1+\nu}$. Since $\mu_{0}$ and $\widetilde{\mu}_{0}$ are both symmetric around origin, we conclude that for any $|i|\leqslant cN$
$|s_{i}(0)-\gamma_{i}(0)|\lesssim\frac{|i|^{1/2}+N^{\nu}}{N}+\frac{i^{2}}{N^{2}}.$
@@
$|s^{\prime}_{i}(0)-\gamma_{i}(0)|\lesssim\frac{|i|^{1/2}+N^{\nu}}{N}+\frac{i^{2}}{N^{2}},\qquad\quad|s_{i}(0)-s^{\prime}_{i}(0)|\lesssim\frac{|i|^{1/2}+N^{\nu}}{N}+\frac{i^{2}}{N^{2}}.$ (4.64)
@@
where $x_{i}(0,\alpha)=\alpha s_{i}+(1-\alpha)s^{\prime}_{i},\quad 0\leqslant\alpha\leqslant 1,$ and the martingales satisfies ${\rm d}\langle b_{i}\rangle/{\rm d}t\leqslant 1$. We consider the corresponding tangential equation (4.20) satisfied by $v_{i}(t)=v_{i}(\alpha,t)=\frac{{\rm d}}{{\rm d}\alpha}x_{i}(\alpha,t)$
@@
${\rm d}f_{t}=m_{t}(z)\partial_{z}f_{t}{\rm d}t+\frac{1}{2N}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}(t)({\rm d}\langle b_{k}\rangle_{t}-{\rm d}t)}{(x_{k}(t)-z)^{3}}-\frac{1}{\sqrt{2N}}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}(t)}{(z-x_{k}(t))^{2}}{\rm d}b_{k}(t).$
@@
${\rm d}f=\sum_{1\leqslant|k|\leqslant N}\frac{{\rm d}v_{k}}{x_{k}-z}+\sum_{1\leqslant|k|\leqslant N}v_{k}{\rm d}\frac{1}{x_{k}-z}.$ (4.67)
[…]
${\rm(I)}$ $=-\frac{1}{\sqrt{2N}}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}}{(z-x_{k})^{2}}{\rm d}b_{k},$
@@
${\rm(III)}$ $=\frac{1}{2N}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}{\rm d}\langle b_{k}\rangle}{(x_{k}-z)^{3}}.$
@@
$\sum_{1\leqslant|k|\leqslant N}\frac{\partial_{t}v_{k}}{x_{k}-z}$ $=\sum_{\ell\neq k}\frac{v_{\ell}-v_{k}}{2N(x_{\ell}-x_{k})^{2}(x_{k}-z)}=\frac{1}{2}\sum_{\ell\neq k}\frac{v_{\ell}-v_{k}}{2N(x_{\ell}-x_{k})^{2}}\left(\frac{1}{x_{k}-z}-\frac{1}{x_{\ell}-z}\right)$
@@
${\rm(II)}+\sum_{1\leqslant|k|\leqslant N}\frac{\partial_{t}v_{k}}{x_{k}-z}=\frac{1}{2N}\sum_{\ell\neq k}\frac{v_{k}}{x_{\ell}-x_{k}}\frac{1}{x_{k}-z}\left(\frac{1}{x_{k}-z}-\frac{1}{x_{\ell}-z}\right)=\frac{1}{2N}\sum_{\ell\neq k}\frac{v_{k}}{(x_{k}-z)^{2}}\frac{1}{x_{\ell}-z}.$
All singularities have disappeared. We obtained ${\rm(II)}+{\rm(III)}+\sum_{1\leqslant|k|\leqslant N}\frac{\partial_{t}v_{k}}{x_{k}-z}=m(z)\partial_{z}f+\frac{1}{2N}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}(t)({\rm d}\langle b_{k}\rangle_{t}-{\rm d}t)}{(x_{k}(t)-z)^{3}}.$ Summation of the remaining term (I) concludes the proof. ∎
@@
For any $|z|\leqslant R/4$ with $E\geqslant 0$, $\eta:=\operatorname{Im}z\geqslant N^{-1+\nu}$, we have
@@
$\begin{split}|\operatorname{Im}f^{(2)}_{0}(z)|&=\left|\operatorname{Im}\sum_{1\leqslant k\leqslant N}\left(\frac{v^{(2)}_{k}(0)}{x_{k}(0)-z}-\frac{v^{(2)}_{k}(0)}{-x_{k}(0)-z}\right)\right|\leqslant\eta\sum_{1\leqslant k\leqslant N}|v^{(2)}_{k}(0)|\left|\frac{1}{|x_{k}-E|^{2}+\eta^{2}}-\frac{1}{|x_{k}(0)+E|^{2}+\eta^{2}}\right|\\ &=4\eta E\sum_{1\leqslant k\leqslant N}|v^{(2)}_{k}(0)|\frac{x_{k}(0)}{|x_{k}(0)-z|^{2}|x_{k}(0)+z|^{2}}\leqslant 4\eta E\sum_{1\leqslant k\leqslant N}\frac{k^{2}}{N^{2}}\frac{x_{k}(0)}{|x_{k}(0)-z|^{2}|x_{k}(0)+z|^{2}}.\end{split}$ (4.68)
From (4.63) and the assumption ${\rm Im}z\geqslant N^{-1+\nu}$ for eigenvalues $x_{k}\leqslant R/2$, and the assumption $x_{k}\leqslant C$ for the complementary regime, we have
@@
$|{\rm Im}f_{0}^{(1)}(z)|\leqslant 4\eta E\sum_{1\leqslant k\leqslant N}\frac{\sqrt{k}}{N}\frac{x_{k}(0)}{|x_{k}(0)-z|^{2}|x_{k}(0)+z|^{2}}\lesssim\sqrt{N}E\int_{0}^{c}\frac{x^{3/2}\eta}{|z-x|^{2}|z+x|^{2}}{\rm d}x\lesssim E\sqrt{\frac{N}{\max(E,\eta)}},$
@@
$\mathscr{S}=\left\{z=E+\mathrm{i}y:y<E<\frac{R}{10}\right\}.$
@@
$\tau_{0}$ $=\inf\left\{0\leqslant t\leqslant 1:\mid\exists z\in[0,R/2]\times[\frac{\varphi^{5}}{N},1]\ {\rm s.t.}\ \left|m_{t}(w)-\widetilde{m}_{t}(w)\right|\geqslant\varphi\sqrt{\frac{{\rm Im}[\widetilde{m}_{0}(w)]}{N{\rm Im}[w]}}\right\},$
$\tau$ $=\min\{\tau_{0},\tau_{\ell,m}:0\leqslant\ell,m\leqslant N^{12}{\mathsf{T}},\varphi^{5}N^{-1}<E_{m}<R/10\},$
@@
$\{\tau={\mathsf{T}}\}\bigcap_{1\leqslant\ell\leqslant N^{12}{\mathsf{T}},1\leqslant m\leqslant N^{12},1\leqslant k\leqslant N}A_{\ell,m,k}\subset\bigcap_{z\in\mathscr{S},0<t<{\mathsf{T}}}\left\{|\operatorname{Im}f_{s}(z_{t-s})-\operatorname{Im}f_{0}(z_{t})|<E\sqrt{\frac{N}{\max\{E,t\}}}+NE\max(E,t)\right\},$ (4.71)